the sides of a triangle are in the ratio 4 5 6. what is the length of each side if the perimeter is 45cm
Answer by ankor@dixie-net.com(12704)

the sides of a triangle are in the ratio 4 5 6. what is the length of each side if the perimeter is 45cm
:
Let x = the multiplier
4x + 5x + 6x = 45
15x = 45
x = 3
;
the 3 sides are:
3*4 = 12
3*5 = 15
3*6 = 18
--------
total 45